OpenStudy (anonymous):

One leg of a right triangle measures 15 m and the hypotenuse measures 17 m. What is the measure of the other leg?

OpenStudy (anonymous):

You can use the same concept as your last question but this time, plug in 17 for c and 15 for a... solve for b to get the length of the leg.

OpenStudy (kirbykirby):

a^2+b^2 = c^2 where c is the hypotenuse. We are given 17 is the hypotenuse. Hence 17^2 = 15^2 + b^2 17^2 - 15^2 = b^2 64 = b^2 b = 8

OpenStudy (kirbykirby):

(at the end, square root both sides)

OpenStudy (anonymous):

When you are asked to find a length of a triangle, you can generally use the formula a^2+b^2=c^2 where a and b are legs and c is the hypotenuse. :)

OpenStudy (anonymous):

thank you